# Bounds on the photon mass via the Shapiro effect in the solar system

### v1

@@ Abstract
We study the effects of a finite mass for the photon on its propagation in a weak gravitational field. In particular, we analyse the gravitational time delay, also known as the Shapiro effect. We work in isotropic coordinates in the weak-field limit and find that the mass-dependent corrections enhance the gravitational time delay. Doppler-tracking data from the Cassini mission allow us to set upper bounds on the photon mass at the level of $m_{\gamma}\lesssim 3\times 10^{-7}\,{\rm eV/c^{2}}$. We also discuss next-generation solar-system tests of general relativity that could improve this upper limit, potentially reaching $m_{\gamma}\lesssim 8\times 10^{-8}\,{\rm eV/c^{2}}$. Though not competitive with the currently best limits, our projected bounds are at the ballpark of earlier ones based on the gravitational deflection angle.
@@ I Introduction
The photon is the paradigm for a massless particle. In fact, the invariance of electromagnetism under local gauge transformations requires $m_{\gamma}=0$, though the photon may indeed acquire an effective mass in certain circumstances Skobelev ; Cover ; supercond . The fundamental masslessness of the photon is a theoretically pleasing hypothesis, usually taken for granted, but it must be thoroughly tested. The most stringent upper bound currently accepted is $m_{\gamma}\leq 10^{-18}\,{\rm eV/c^{2}}\approx 10^{-51}$ g extracted from data from the Voyager missions Ryutov ; for reviews, see the latest edition of the Particle Data Group PDG , as well as refs. Tu ; Nieto ; Okun ; Goldhaber ; Goldhaber2 .
@@
Here we study the effect of the photon mass on the gravitational time delay, first suggested by Shapiro in 1964 as a “fourth test of general relativity” Shapiro1 . His idea was to send radio signals from radio telescopes on Earth towards Mercury or Venus when these are close to superior conjunction, i.e., passing almost behind the Sun. According to Shapiro’s calculations, the round-trip time of the signal would be longer than expected in the absence of the Sun: for grazing incidence at the solar limb, the extra travel time due to gravity can reach $\sim 200\,{\rm\mu s}$ in the inner solar system.
@@
The measurements were conducted by Shapiro et al. Shapiro2 ; Shapiro3 using the Arecibo and Haystack telescopes with dedicated radio systems, which allowed the detection of the extremely faint echoes ($\approx 10^{27}$ times weaker than the emitted signals). Curvature effects in general relativity are characterized by Eddington’s $\gamma$ parameter, which is unity in general relativity Poisson ; Will_book , and the data analysis yielded $\gamma=1.03\pm 0.04$, thus confirming the predictions of general relativity to within a few percent. Other similar analyses were undertaken using ranging data from the Mariner landers Mariner1 and the Viking spacecraft Viking ; Viking2 ; Viking3 , both sent to Mars, determining $\gamma$ to within $3\%$ and $0.2\%$ accuracy, respectively. The most precise determination of $\gamma$ so far stems from Doppler-tracking data from the Cassini spacecraft on its way to Saturn, which yielded $|\gamma-1|=2.3\times 10^{-5}$ Cassini_1 ; Cassini_2 . These results will be explored in Sec. III.3 to set upper bounds on the photon mass.
@@
Quite generally, the search for deviations of general relativity, or new-physics effects in general, can profit greatly from experiments in space Book_space . Tests of the universality of free fall, for example, may be performed over longer time periods, thus enabling higher precision levels to be reached. Similarly, interferometric experiments can benefit from much longer interaction times, as well as longer distances, both helping to enhance possible signals. The absence of seismic vibrations and atmospheric disturbances, such as optical distortions and drag (which imparts non-gravitational accelerations to sensors) are further advantages of space-based experiments. Also for time-delay measurements the space environment is ideal, since precise time keeping is crucial when minute height and time inaccuracies may impact the results. In space these variables can be particularly well controlled, thus allowing for very precise measurements.
@@
So far all measurements of the Shapiro effect were performed using ground-based antennas sending and receiving signals to spacecraft at interplanetary distances. Now, however, the technologies necessary to measure time delays to picosecond accuracy in space are mature (e.g., highly stable laser sources, ultra-precise atomic clocks, drag-free control, etc Book_space ). In this context, we also analyse a few promising space missions which include time-delay measurements as their scientific targets, aiming at drastically improving the current limits on $|\gamma-1|$. The projected uncertainties in $\gamma$ will be used to derive sensitivities to the photon mass, cf. Sec. III.4.
@@
This paper is organized as follows: in Sec. II we set up the theoretical basis of our analysis and find the gravitational time delay in the case of massive photons. Next, in Sec. III we apply our result to current and future limits on $|\gamma-1|$ to extract upper bounds on the photon mass. Finally, in Sec. IV we present our conclusions. We use SI units throughout.
@@ II Gravitational effects
We are interested in the behavior of massive Abelian spin-1 bosons propagating in a weak gravitational field. The dynamics of the massive photon is governed by the de Broglie-Proca action coupled to gravity dB1 ; dB2 ; dB3 ; Proca1 ; Proca2
@@
Before starting we must decide on the coordinates to describe the spacetime geometry of a point mass. While standard Schwarzschild coordinates are common, here we employ Sun-centered isotropic spherical coordinates with which the interval squared becomes
@@
Isotropic coordinates are typically used in astrometric calculations and are particularly suited for comparison with observational data. Also important in our case is the fact that the post-Newtonian terms in the context of a parametrized post-Newtonian (PPN) expansion Will_book ; Will_exp ; Turyshev are expressed in these coordinates. Our choice therefore simplifies the job of comparing our results including a finite photon mass to the PPN expressions containing Eddinton’s $\gamma$ parameter, cf. App. A.
In Sec. II.1, as a warm up, we derive the deflection angle for a massive photon passing close to the Sun. Next, in Sec. II.2, we determine the gravitational time delay. In order to capture both gravitational and curvature effects, we will use the actual geodesic equations, obtained in App. B for convenience.
@@ II.1 Deflection angle
One of the three classic tests of Einstein’s general relativity is the deflection of light by a massive object, which turns out to be twice as large as that derived from Newtonian mechanics: $\delta_{\rm E}=2\delta_{\rm N}=4GM_{\odot}/c^{2}r_{0}$, where $r_{0}$ is the distance of closest approach (for the Sun, it is the perihelion). Light, typically taken to be exactly massless, is just an extreme test case of the theory, since massive particles are obviously also affected by curvature. We are interested in the latter case, where photons are exceedingly light, but massive nevertheless.
In App. B we derived the geodesic equations and showed that we are allowed to constrain the trajectory to the equatorial plane ($\theta=\pi/2$) without loss of generality. Our target is to obtain an expression for $d\phi/dr$ as a function of the radial coordinate and, to this end, we start with eq. (42d) writing it as
@@
and $\alpha=m_{\gamma}c^{2}/E$. Plugging eq. (3) into the constraint $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=c^{2}$, cf. eq. (42c), we find
@@
Equation (9) matches the results from Refs. Accioly1 ; Accioly2 calculated via scattering methods, as well as those from a simplified calculation using a straight-line approximation, cf. App. A. In Ref. Accioly2 the authors use 43-GHz data from the Very-Long Baseline Array (VLBA) Fomalont and derive $m_{\gamma}\lesssim 3.5\times 10^{-5}\,{\rm eV/c^{2}}$, whereas in Ref. Accioly1 interferometric techniques are also used, but with 2-GHz radio waves Lebach , to obtain [^2] In Ref. Accioly1 the authors quote $m_{\gamma}\lesssim 10^{-40}\,{\rm g}\approx 5.6\times 10^{-8}\,{\rm eV/c^{2}}$. However, the deviation of the measured angles relative to Einstein’s result is $\Delta/\delta_{\rm E}\approx 8\times 10^{-4}$ Lebach , so that using the first equation in their section 3 we get $m_{\gamma}\lesssim 5.9\times 10^{-40}\,{\rm g}\approx 3.3\times 10^{-7}\,{\rm eV/c^{2}}$, a factor six worse than their stated limit. $m_{\gamma}\lesssim 3.3\times 10^{-7}\,{\rm eV/c^{2}}$. In Sec. III we will extract bounds complementary to those.
@@ II.2 Gravitational time delay
Our objective is to determine the total (coordinate) time elapsed from emission of the signal on Earth until its arrival at the receiver. Therefore, we must find $dt/dr$ as a function of known parameters, in particular the heliocentric radial distances of the objects involved. Once again we constrain the movement to the equatorial plane. We do not need to go back to the geodesic equations: it suffices to divide the constraint equation (42c) by $\dot{r}$ to get
@@
which gives a constant correction factor deviating from unity by [^3] The mean orbital radius of the Earth is $r_{e}=149\times 10^{6}$ km. $\approx 10^{-8}$.
With eqs. (II.2) and (16) the proper time elapsed between emission and arrival of the signal is finally
@@
where we neglected terms of order $\mathcal{O}(\alpha^{2}GM_{\odot}/c^{2}r_{e})$, as well as post-post-Newtonian corrections. It is worth keeping in mind that the gravitational contributions are of order $GM_{\odot}/c^{3}\approx 5\,{\rm\mu s}$. Note that finite-mass corrections affect only the first and third terms, but crucially not the phenomenologically most relevant logarithmic term. This will have important consequences when we extract limits on the photon mass in Sec. III.
@@
Before closing this section, let us make a few general comments about our main result, eq. (II.2). The first term represents the Newtonian travel time in the absence of the Sun. This is effectively an Euclidean distance which cannot be measured directly due to the actual curvature of spacetime. The mass dependence of this term is expected, since the time it takes for a particle with normalized velocity $\beta=v/c$ to cover a distance $\ell$ is $(\ell/c)\beta^{-1}$. But in special relativity $E/m_{\gamma}c^{2}=(1-\beta^{2})^{-1/2}=\alpha^{-1}$, which can be rearranged as $\beta^{-1}=(1-\alpha^{2})^{-1/2}$. Expanding this expression yields $\beta^{-1}\approx 1+\alpha^{2}/2$, thus matching the mass-dependent factor in the first line of eq. (II.2). The other terms are general relativistic in nature and represent an extra travel time due to gravity. The gravitational time delay, or Shapiro effect, is defined as the non-Newtonian contribution to the total round-trip time: $\Delta\tau_{\rm delay}=\Delta\tau-\Delta\tau(M_{\odot}=0)$.
It is worth noting that in his first work Shapiro1 , Shapiro used Schwarzschild coordinates to compute the (proper) time delay in the straight-line approximation, but in his final analysis in Ref. Shapiro3 he mentions the use of harmonic coordinates and takes only the logarithm term as representative of the gravitational time delay. This is sensible, since this term is the most sensitive to the perihelion distance, $r_{0}$, which is the only time-dependent variable. Moreover, the logarithm term is the same in the simplified straight-line approximation, cf. App. A.1.
III Experimental sensitivities
@@
In Sec. II.2 we obtained the correction to the two-way travel time of a signal as measured by an observer on Earth, mediated by massive photons, to first post-Newtonian order. Now we would like to use this result, together with current and future limits on deviations from standard general relativity (where the photon is massless), to set upper bounds on a putative rest mass of the photon.
Generally, the Shapiro effect refers to the non-Newtonian contribution to the round-trip time difference. In the context of tests of general relativity the gravitational time delay is expressed in the parameterized post-Newtonian (PPN) metric in terms of Eddington’s $\gamma$ parameter Poisson ; Will_book ; Will_exp , which characterizes curvature effects. Incidentally, it is also relevant in analyses of the deflection angle, cf. Sec. II.1.
As mentioned at the end of Sec. II.2, the logarithmic term is characteristic of the Shapiro effect due to its stronger dependence on the perihelion distance. Accordingly, the PPN expression for the time delay is usually quoted in the form of eq. (31), which was derived using a straight-line approximation for the light path (see discussion in App. A). However, using the actual geodesic we get eq. (37) and applying the correction factor (16) we find
$\Delta\tau_{\rm delay}^{\rm PPN}$ $=$ $\frac{4GM_{\odot}}{c^{3}}\left(\frac{1+\gamma}{2}\right)\left[2+\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)\right]$ (18)
-- $\frac{3GM_{\odot}}{c^{3}}\left(1+\frac{r_{r}}{r_{e}}\right)\,,$
where we assumed that $r_{0}\ll r_{e},r_{r}$, which is valid in all cases analysed here. Clearly, the logarithmic term is not the only contribution at post-Newtonian order. This expression will be our standard for comparisons with measurements of Eddington’s $\gamma$ parameter.
III.1 Bound from the Mariner 6 and 7 missions
The ten Mariner spacecraft were part of a program to study the inner solar system. In particular, Mariner 6 and 7 completed their flyby of Mars in 1969, whereupon they analysed its atmosphere and surface, relaying data back to Earth using 2.2 GHz (S-band) radio signals. Soon thereafter, the spacecraft entered heliocentric orbits.
In what follows we assume that the signals from emitter to receiver and back have approximately the same frequency. Subtracting the (unmeasurable) Newtonian light time from eq. (II.2) and comparing with eq. (18) we are able to express the deviation from general relativity, parametrized by $|\gamma-1|$, in terms of the photon mass as
$|\gamma-1|=\frac{4\alpha^{2}}{2+\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)}\,.$ (19)
Shorter perihelion distances increase the denominator, thereby reducing $|\gamma-1|$. This is sensible, since in this case light rays experience stronger gravitational fields and deviations from general relativity may be probed to better accuracy. On the other hand, this also means that our bounds will be weaker at shorter perihelion distances, a direct consequence of the fact that the logarithmic term is not affected by the finite photon mass, cf. eq. (II.2).
At this point it is useful to write $\alpha=m_{\gamma}c^{2}/E$ more explicitly as
where we used the quantum-mechanical relation $E=hf$. Assuming that no clear deviation from general relativity is found, any signs of new physics must lie below the uncertainty in $\gamma$, which is encoded in $|\gamma-1|$. Thus, using eq. (20) we have the following sensitivity
$m_{\gamma}$ $\lesssim$ $2.1\times 10^{-9}\,{\rm eV/c^{2}}\left(\frac{|\gamma-1|}{10^{-6}}\right)^{1/2}\left(\frac{f}{1\,{\rm GHz}}\right)$ (21)
$\times$ $\sqrt{2+\log\left(\frac{4r_{e}r_{r}}{r_{0}^{2}}\right)}\,.$
The behavior of eq. (21) as a function of the perihelion distance $r_{0}$ is displayed in Fig. 1.
Figure 1: The sensitivity (21) for $r_{e}=r_{r}=1$ AU, $|\gamma-1|=10^{-6}$ and $f=1$ GHz as a function of the perihelion distance $r_{0}$ in units of the solar radius. The slope is typical of the Shapiro time delay, cf eq. (II.2).
According to Ref. Mariner1 , superior solar conjunction for Mariner 6 and 7 happened around mid 1970 when the Earth-spacecraft distances were $402\times 10^{6}$ km (Mariner 6) and $386\times 10^{6}$ km (Mariner 7) with Sun-Earth-spacecraft angles of $0.95^{\circ}$ and $1.56^{\circ}$, respectively. This implies [^4] We use the Euclidean law of cosines as a good approximation. that $r_{r}\approx 253\times 10^{6}$ km and $r_{0}\approx 3.6R_{\odot}$ for Mariner 6, and $r_{r}\approx 237\times 10^{6}$ km and $r_{0}\approx 5.9R_{\odot}$ for Mariner 7. With the quoted uncertainty of $3\%$ on $\gamma$, we may use eq. (21) to calculate the upper bounds on the photon mass based on the data from each spacecraft. As it turns out, the results are very similar, differing only by $\approx 5\%$, so we take the weakest (from Mariner 6) as our conservative estimate, namely
$m_{\gamma}\lesssim 2.8\times 10^{-6}\,{\rm eV/c^{2}}\,.$ (22)
III.2 Bound from the Viking 1 and 2 missions
The Viking 1 and 2 missions, launched in 1975, had the detailed study of the Martian surface and atmosphere, and the search for signs of life as their main scientific goals. During their operation (at $r_{r}=1.52$ AU), the spacecraft worked with radio signals at $2.2$ GHz (S-band) with impact parameters as small as $r_{0}\approx 1.5R_{\odot}$ close to superior conjunction in late 1976 Viking ; Viking2 . The attained precision of $\approx 10$ ns led to a determination of $|\gamma-1|$ to within $0.2\%$ Viking3 . With this, eq. (21) gives us
$m_{\gamma}\lesssim 7.7\times 10^{-7}\,{\rm eV/c^{2}}\,,$ (23)
slightly improving the bounds from Mariner 6 and 7.
III.3 Bounds from the Cassini-Huygens mission

### v2

@@ Abstract
We study the effects of a finite mass for the photon on its propagation in a weak gravitational field. In particular, we analyse the gravitational time delay, also known as the Shapiro effect. We work in isotropic coordinates in the weak-field limit and find that the mass-dependent corrections enhance the gravitational time delay. Doppler-tracking data from the Cassini mission allow us to set an upper bound on the photon mass, namely $m_{\gamma}<4.9\times 10^{-7}\,{\rm eV/c^{2}}$ at $95\%$ CL. We also discuss next-generation solar-system tests of general relativity that could improve this upper limit, potentially by a factor of ten. Though not competitive with the currently best limits, our bound is at the ballpark of earlier ones based on the gravitational bending of light by the Sun.
@@ I Introduction
The photon is the paradigm for a massless particle. In fact, the invariance of electromagnetism under local gauge transformations requires $m_{\gamma}=0$, though the photon may indeed acquire an effective mass in certain circumstances Skobelev ; Cover ; supercond . The fundamental masslessness of the photon is a theoretically pleasing hypothesis, usually taken for granted, but it must be thoroughly tested. The most stringent upper bound currently accepted is $m_{\gamma}<10^{-18}\,{\rm eV/c^{2}}$ extracted from data from the Voyager missions Ryutov ; for reviews, see the latest edition of the Particle Data Group PDG , as well as refs. Tu ; Nieto ; Okun ; Goldhaber ; Goldhaber2 .
@@
Here we study the effect of the photon mass on the gravitational time delay, first suggested by Shapiro in 1964 as a “fourth test of general relativity” Shapiro1 . His idea was to send signals from radio telescopes on Earth towards Mercury or Venus when these are close to superior conjunction, i.e., passing almost behind the Sun. According to Shapiro’s calculations, the round-trip time of the signal would be longer than expected in the absence of the Sun: for grazing incidence at the solar limb, the extra travel time due to gravity can reach $\sim 200\,{\rm\mu s}$ in the inner solar system.
@@
The measurements were conducted using the Arecibo and Haystack telescopes with dedicated radio systems allowing the detection of the faint echoes ($\approx 10^{27}$ weaker than the emitted signals) Shapiro2 ; Shapiro3 . Curvature effects are characterized by Eddington’s $\gamma$ parameter, which is unity in standard general relativity Poisson ; Will_book , and the data analysis yielded $\gamma=1.03\pm 0.04$, thus confirming the predictions of general relativity to within a few percent. Other similar analyses were undertaken using ranging data from the Mariner landers Mariner1 and the Viking spacecraft Viking ; Viking2 ; Viking3 , both sent to Mars, determining $\gamma$ to within $3\%$ and $0.2\%$ accuracy, respectively. The most precise determination of $\gamma$ so far stems from Doppler-tracking data from the Cassini spacecraft on its way to Saturn, which yielded $|\gamma-1|=2.3\times 10^{-5}$ Cassini_1 ; Cassini_2 . These results will be explored in Sec. III to constrain the photon mass.
@@
Quite generally, the search for deviations of general relativity, or new-physics effects in general, can profit greatly from experiments in space Book_space . Tests of the universality of free fall, for example, may be performed over longer time periods, thus reaching higher precision levels. Similarly, interferometric experiments can benefit from much longer interaction times, as well as longer distances, both helping to enhance possible signals. The absence of seismic vibrations and atmospheric disturbances, which impart non-gravitational accelerations to sensors, are further advantages of space-based experiments. Also for time-delay measurements the space environment is ideal, since precise time keeping is crucial when minute height and time inaccuracies may impact the results. In space these variables can be particularly well controlled, thus allowing for very precise measurements.
@@
So far all measurements of the Shapiro effect were performed using ground-based antennas sending and receiving signals to spacecraft at interplanetary distances. Now, however, the technology necessary to measure time delays to picosecond accuracy in space is mature (e.g., with highly stable laser sources, ultra-precise atomic clocks, drag-free control, etc Book_space ). In this context, we also analyse a couple of promising space missions which include time-delay measurements as their scientific targets aiming at considerably improving the current limit on $|\gamma-1|$.
@@
This paper is organized as follows: in Sec. II we set up the theoretical basis of our analysis and find the gravitational time delay in the case of massive photons. Next, in Sec. III we use data from the Cassini time-delay experiment to extract an upper bound on the photon mass. Finally, in Sec. IV we present our conclusions and briefly comment on future prospects from upcoming space-based missions. We use SI units throughout.
@@ II Gravitational effects
We are interested in the behavior of a heavy photon – a massive and electrically neutral Abelian spin-1 boson – propagating in a weak gravitational field. The dynamics of this particle is governed by the de Broglie-Proca action coupled to gravity dB1 ; dB2 ; dB3 ; Proca1 ; Proca2 :
@@
Before starting we must decide on the coordinates to describe the spacetime curvature caused by a central mass (here we neglect the its finite size). While standard Schwarzschild coordinates are common, here we employ heliocentric isotropic spherical coordinates with which the interval squared is given by
@@
Isotropic coordinates are typically used in astrometric calculations and are particularly suited for comparison with observational data. Also important in our case is the fact that the post-Newtonian terms in the context of a parametrized post-Newtonian (PPN) expansion Will_book ; Will_exp ; Turyshev are expressed in these coordinates. Our choice therefore simplifies the comparison between our results including a finite photon mass and the PPN expressions containing Eddinton’s $\gamma$ parameter, cf. App. A.
@@ II.1 Deflection angle
One of the three classic tests of Einstein’s general relativity is the deflection of light by a massive object, which is twice as large as the Newtonian prediction: $\delta_{\rm E}=2\delta_{\rm N}=4GM_{\odot}/c^{2}r_{0}$, where $r_{0}$ is the distance of closest approach (for the Sun it is the perihelion). Light, typically taken to be exactly massless, is just an extreme test case of the theory, since massive particles are obviously also affected by curvature. We are interested in the latter case, where photons are exceedingly light, but massive nevertheless.
In App. B we derived the geodesic equations [^2] It may be shown that, in the limit of geometric optics Harte_2018 , the inclusion of a mass scale does not modify the fact that the constant phase surface is normal to the wave vector, which satisfies a geodesic equation. and showed that we are allowed to constrain the trajectory to the equatorial plane ($\theta=\pi/2$) without loss of generality. Our target is to obtain an expression for $d\phi/dr$ as a function of the radial coordinate and, to this end, we start with eq. (44d) writing it as
@@
and $\alpha=m_{\gamma}c^{2}/E$, which may be expressed as
where we used the quantum-mechanical relation $E=hf$. Plugging eq. (3) into the constraint $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=c^{2}$, cf. eq. (44c), we find
@@
Equation (10) matches the results from Refs. Accioly1 ; Accioly2 calculated via scattering methods, as well as those from a simplified calculation using a straight-line approximation, cf. App. A. In Ref. Accioly2 the authors use 43-GHz data from the Very-Long Baseline Array (VLBA) Fomalont and derive $m_{\gamma}\lesssim 3.5\times 10^{-5}\,{\rm eV/c^{2}}$, whereas in Ref. Accioly1 interferometric techniques are also used, but with 2-GHz radio waves Lebach , to obtain [^3] In Ref. Accioly1 the authors quote $m_{\gamma}\lesssim 10^{-40}\,{\rm g}\approx 5.6\times 10^{-8}\,{\rm eV/c^{2}}$. However, the deviation of the measured angles relative to Einstein’s result is $\Delta/\delta_{\rm E}\approx 8\times 10^{-4}$ Lebach , so that using the first equation in their section 3 we get $m_{\gamma}\lesssim 5.9\times 10^{-40}\,{\rm g}\approx 3.3\times 10^{-7}\,{\rm eV/c^{2}}$, a factor six worse than their stated limit. $m_{\gamma}\lesssim 3.3\times 10^{-7}\,{\rm eV/c^{2}}$.
@@ II.2 Gravitational time delay
We want to determine the total (coordinate) time elapsed from emission of the signal on Earth until its arrival at the receiver, so we must find $dt/dr$ as a function of known parameters, in particular the heliocentric radial distances of the objects involved. Once again we constrain the movement to the equatorial plane. We do not need to go back to the geodesic equations: it suffices to divide the constraint equation (44c) by $\dot{r}$ to get
@@
[…]
With eqs. (II.2) and (17) the proper time elapsed between emission and arrival of the signal on Earth (round trip) is finally
@@
where we neglected terms of order $\mathcal{O}(\alpha^{2}GM_{\odot}/c^{2}r_{e})$, as well as post-post-Newtonian corrections. It is worth keeping in mind that the gravitational contributions are of order $GM_{\odot}/c^{3}\approx 5\,{\rm\mu s}$. Note that finite-mass corrections affect only the first and third terms, but crucially not the phenomenologically most relevant logarithmic term.
@@
Before closing this section, let us make a few general comments about our main result, eq. (II.2). The first term represents the Newtonian travel time in the absence of the Sun. This is effectively an Euclidean distance which cannot be measured directly due to the actual curvature of spacetime. The mass dependence of this term is expected, since the time it takes for a particle with normalized velocity $\beta=v/c$ to cover a distance $\ell$ is $(\ell/c)\beta^{-1}$. But in special relativity $E/m_{\gamma}c^{2}=(1-\beta^{2})^{-1/2}=\alpha^{-1}$, which can be rearranged as $\beta^{-1}=(1-\alpha^{2})^{-1/2}$. Expanding this expression yields $\beta^{-1}\approx 1+\alpha^{2}/2$, thus matching the mass-dependent factor in the first line of eq. (II.2). The other terms are general relativistic in nature and represent an extra travel time due to gravity. The gravitational time delay, the so-called Shapiro effect, is defined as the non-Newtonian contribution to the total round-trip time,
$\Delta\tau_{\rm delay}=\Delta\tau-\Delta\tau(M_{\odot}=0)\,.$ (19)
From now on we shall consider only the time delay as defined above.
III Bound on the photon mass via Cassini data
@@
In Sec. II.2 we obtained the correction to the round-trip travel time of a signal as measured by an observer on Earth, mediated by massive photons, to first post-Newtonian order, cf. eq. (II.2). In order to constrain a putative rest mass of the photon we must compare our theoretical model to observational data.